OpenStudy (anonymous):

For the function, find the y intercept. y=-5/6(x+12)^2

OpenStudy (anonymous):

To find the intercept with Y, just match x to 0, so y=-5/6(0+12)^2=-5/864

OpenStudy (anonymous):

To fin intercepts with X, just solve the equation y=0

OpenStudy (anonymous):

so to find the x intercept, y=0. 0=-5/6... help.

OpenStudy (anonymous):

Just one question, y is \[-\frac{5}{6(x+12)^2}\] or \[-\frac{5}{6}(x+12)^2\] ?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

help me find the x :c

OpenStudy (anonymous):

if \[y=-\frac{5}{6}(x+12)^2\] Y-axis intercept \[y=-\frac{5}{6}(0+12)^2=-5\] x-axis incercept \[0=-\frac{5}{6}(x+12)^2\\0=(x+12)^2\\x=-12\]

OpenStudy (anonymous):

This is the graphic of y: